# Positive paths in Diffeomorphism groups of manifolds with a contact distribution

### v1

@@ 1. Introduction
Let $(M,\xi)$ be a (not necessarily closed) cooriented contact manifold of dimension $12n+$ 1 for $n\geq 1$. There are two natural groups associated with $M$: The group of diffeomorphisms $\operatorname*{Diff}(M)$ and the group of contactomorphisms $\operatorname*{Cont}(M,\xi)$, i.e., the subgroup of diffeomorphisms preserving the contact structure $\xi$.
@@ 1.1. Positivity for contactomorphisms and diffeomorphisms
In [14] the authors introduce the notion of a positive path of contactomorphisms (see also [6]) as follows. Any smooth path $(\varphi_{t})_{t\in I}$ in $\operatorname*{Cont}(M,\xi)$ defines a time-dependent contact vector field $(X_{t}^{\varphi})_{t\in I}$ on $M$. The path is called positive (non-negative) if $\alpha\left(X_{t}^{\varphi}\right)>0$ ($\geq 0$) for all $t$, where $\alpha$ is any contact form defining the chosen coorientation of $\xi$. As observed in [14] for the case of $\widetilde{\operatorname*{Cont}(M,\xi)}$ and [10, 11, 8] for $\operatorname*{Cont}(M,\xi)$, the notion of positivity can be used to define a bi-invariant relation on $\operatorname*{Cont}(M,\xi)$ (and its universal cover $\widetilde{\operatorname*{Cont}(M,\xi)}$). This relation turns out to be a partial order if and only if there are no positive (contractible) loops of contactomorphisms [14].
The notion of positivity has turned out to be connected to basically all known phenomena in contact geometry, such as for example the existence of closed Reeb orbits [2], contact (non-)squeezing [13, 4], the existence of invariant norms [12, 16], translated points [27, 2, 19], (non-) displaceability of pre-Lagrangians [7, 23], loose Legendrians [22] or supporting open book decompositions [19].
In principle the same notion of positivity can be defined for paths in $\operatorname*{Diff}(M)$, see [15, Remark 5.4]. Since any path $(f_{t})_{t\in I}$ defines a time-dependent vector field $(X_{t}^{f})_{t\in I}$ on $M$, we can analogously define a path of diffeomorphisms to be positive (non-negative) if $\alpha\left(X_{t}^{f}\right)>0$ ($\geq 0$) for all $t$. Note however that this notion of positivity is in general not invariant under conjugation. Another major difference to the contact case is that there exist non-constant paths of diffeomorphisms with $\alpha\left(X_{t}^{f}\right)=0$. We call a path $(f_{t})_{t\in I}$ with this property a null path.
In this paper we show that the notion of positivity is trivial on the identity component of the diffeomorphism group $\operatorname*{Diff}_{0}(M)$ in the following sense.
@@
Consider $\mathbb{R}^{2n+1}$ equipped with its standard contact structure. Every $f\in\operatorname*{Diff}_{0}(\mathbb{R}^{2n+1})$ is connected to $\mathrm{id}$ by a null path and a positive path of diffeomorphisms. If $f$ can be connected to $\mathrm{id}$ by a path of compactly supported diffeomorphisms, it is connected to $\mathrm{id}$ by a compactly supported null path and a positive path that is arbitrarily close to $\mathrm{id}$ in the $C^{\infty}$-compact open topology outside of a compact set.
@@
The following shows that it is possible to extend a given path of contactomorphisms that is positive outside a compact subset to an everywhere positive path of diffeomorphisms with the same endpoints.
@@
Consider $\mathbb{R}^{2n+1}$ equipped with its standard contact structure. Let $(f_{t})_{t\in[0,1]}$ be a path of contactomorphisms such that on the complement of some compact subset $K_{1}$ the path is positive. Then there exists a positive path of diffeomorphisms $(g_{t})_{t\in[0,1]}$ connecting $f_{0}$ and $f_{1}$ such that $g_{t}$ coincides with $f_{t}$ on the complement of some compact subset $K_{2}\supset K_{1}$.
For general contact manifolds one can show.
@@
For the case of compactly supported diffeomorphisms the local contractibility of the $C^{\infty}$-topology implies.
@@
Even if a positive loop of contactomorphisms is contractible, it is in general not possible to contract it in $\operatorname*{Cont}_{0}(M,\xi)$ via non-negative loops [13, Theorem 1.11]. This changes for diffeomorphisms.
@@ 1.2. Positive paths between submanifolds and thermodynamic processes
As pointed out in [15, Remark 5.4] one can also define a notion of positivity for smooth families of embeddings into a contact manifold. Let $N$ be a smooth manifold and $\iota\colon N\times[0,1]\hookrightarrow M$ be a smooth family of embeddings. We assume that all embeddings are proper in the sense that the intersection of the image with any compact subset in $M$ is compact. Call $\iota$ positive/ non-negative/ null if $\alpha(d\iota(\partial_{t}))$ is positive/ non-negative/ null, i.e., if the path $t\mapsto\iota(p,t)$ is positively/non-negatively transverse or tangent to $\xi$ for any $p\in N$.
@@
In [15] the authors look at positive (non-negative) isotopies of Legendrians in the context of thermodynamic processes. More precisely, they look at sets of equilibrium states in a thermodynamic phase space $\mathcal{T}$. The space $\mathcal{T}$ naturally carries a contact structure and is contactomorphic to the $1$-jet bundle $J^{1}\mathbb{R}^{n}\cong\mathbb{R}^{2n+1}$ of $\mathbb{R}^{n}$ (or more generally of any smooth manifold $X$) equipped with its canonical contact structure. Sets of equilibrium states naturally form Legendrian submanifolds in this space, i.e., submanifolds $L$ with $TL\subset\xi$. As before we assume that $\mathcal{T}$ can be covered by finitely many Darboux charts. Slow thermodynamic processes can be described in this language by compactly supported non-negative isotopies of Legendrians, see [15, Section 5]. In [15, Remark 5.4] the authors consider smooth (possibly non-Legendrian) isotopies between two Legendrians and raise the question if the non-negativity assumption is restrictive in this setting. The following corollary of Theorem 1.1 and Theorem 1.3 provides a negative answer.
@@
Let $L_{1},L_{2}$ be Legendrians in $\mathcal{T}$ that are isotopic via a compactly supported smooth isotopy. Then there exist a positive and a compactly supported null isotopy between $L_{1}$ and $L_{2}$. If $\mathcal{T}\cong J^{1}\mathbb{R}^{n}$ the statement is true for any two Legendrians $L_{1},L_{2}$ such that there exist $f\in\operatorname*{Diff}_{0}(J^{1}\mathbb{R}^{n})$ with $f(L_{1})=L_{2}$. Moreover, in this case the positive isotopy can be chosen $C^{\infty}$-compact open close to the constant isotopy in the complement of a compact subset.
@@ 1.3. Norms on groups of diffeomorphisms
A further direct consequence of Theorem 1.3 and Corollary 1.6 is that, different from contact geometry [28, 26], the notion of contact Hamiltonian can not be used to define any meaningful Hofer-type (psudo-) norms or metrics on the group $\operatorname*{Diff}_{0}(M)$ or the isotopy class of any submanifold.
For instance given a path of diffeomorphisms $(f_{t})_{t\in[0,1]}$ on a closed manifold generating time-dependent vector field $X_{t}$, one can define its Hofer length with respect to a contact form $\alpha$ as
@@ Acknowledgements
I thank Leonid Polterovich for his comments on a first version of the paper and for encouraging to prove parts of Theorem 1.1 and Theorem 1.2. I also thank Egor Shelukhin for valuable feedback. This work was supported by the CRM-ISM and CIRGET postdoctoral fellowships and by the Fondation Courtois.
@@ 2.1. Notations
Throughout this note $(M,\xi)$ will be a (not necessarily closed) contact manifold of dimension $12n+$ 1 for some $n\geq 1$. We assume that $(M,\xi)$ is cooriented, i.e., that $\xi$ is the kernel of some global $1$-form $\alpha$ such that $\alpha\wedge d\alpha^{n}$ is a volume form on $M$.
@@
A smooth path of diffeomorphisms $(\varphi_{t})_{t\in I}$ is a family of diffeomorphisms which is smooth as a map $I\times M\rightarrow M$. Given a possibly time dependent vector field $X_{t}$, we will denote its flow at time $t$ by $\varphi_{t}^{X}$ and the time dependent vector field generated by a path of diffeomorphisms $\varphi_{t}$ by $X_{t}^{\varphi}$.
@@ 3.1. Diffeomorphisms of Darboux-balls
The following theorem is a variation of Proposition 2.2 for the case of the standard contact structure. The main difference here is that for the standard contact structure the above vector fields used in Proposition 2.2 can be chosen linearly independent on all of $\mathbb{R}^{2n+1}$. This allows to extend the result to non-compactly supported diffeomorphisms.
@@
(1) $f=\varphi^{X}_{\tau_{1}\circ\Phi_{2}^{-1}}\circ\varphi^{Y}_{\tau_{2}\circ\Phi_{1}^{-1}}\circ\varphi^{Z}_{\tau_{3}}=\varphi^{X}_{\tau_{1}\circ\Phi_{2}^{-1}}\circ\varphi^{Y}_{\tau_{2}\circ\Phi_{1}^{-1}}\circ\varphi^{X}_{\epsilon}\circ\varphi^{Y}_{\tau_{3}\circ\varphi_{\epsilon}^{X}}\circ(\varphi^{X}_{\epsilon})^{-1}.$
@@
Let $\varphi_{t}$ be the Reeb-flow of some contact form $\alpha$, i.e., $\varphi_{t}$ is the flow of a vector field $R$ uniquely defined by the equations $\alpha(R)=1$ and $d\alpha(R,\cdot)=0$. Then $\varphi_{t}$ is a positive path of diffeomorphisms preserving the contact form $\alpha$. Pick a null path of diffeomorphisms $(g_{t})_{t\in[0,1]}$ with $g_{0}=\mathrm{id}$ and $g_{1}=\varphi_{-1}f$. The composition $(\varphi_{t}g_{t})_{t_{i}n[0,1]}$ is a positive path of diffeomorphisms from $\mathrm{id}$ to $f$.
@@
(4) $\varphi^{Y}_{s(\tau_{3}\circ\varphi_{\epsilon}^{X})}$ $=\left(x,y+\frac{st}{\epsilon},z+(\epsilon-x)\frac{st}{\epsilon}\right)$
@@
Let $\mathcal{U}_{co}$ be a neighbourhood of the identity in the $C^{\infty}$-compact open topology. For $|t|$ sufficiently small the Reeb flow $\varphi_{t}$ at time $t$ lies in $\mathcal{U}_{co}$ and from the above expressions one sees that, picking $|t|$ small enough, $\varphi_{t}$ is connected to $\mathrm{id}$ by a null path contained in $\mathcal{U}_{co}$. Given a compactly supported path $(f_{s})_{s\in[0,1]}$ and fixed $t$ small enough one gets a null path $g_{s}$ connecting $\mathrm{id}$ and $\varphi_{-t}\circ f_{1}$ that is arbitrarily close in the compact open topology to the identity outside of some compact subset. As before the desired positive path we get by looking at the composition $\varphi_{st}\circ g_{s}$.
@@
Let $(f_{t})_{t\in[0,1]}$ be a path of contactomorphisms that is positive outside a compact set $K_{0}$. We can without restriction assume that $f_{0}=\mathrm{id}$. Pick a compact set $K_{1}$ such that $f^{-1}_{t}(K_{0})\subset K_{1}$ for all $t\in[0,1]$. Given a contact form $\alpha$, let $C_{1}$ be the minimum of the time-dependent contact Hamiltonian inducing the path $(f_{t})_{t\in[0,1]}$. Note that $f_{t}^{\ast}\alpha=\rho_{t}\alpha$ for some family of positive functions $(\rho_{t})_{t\in[0,1]}$. Let $C_{2}$ be the maximum of the family $(\rho_{t})_{t\in[0,1]}$ on $K_{1}$. Pick compact subsets $K_{2},K_{3}$ with $K_{1}\subset K_{2}\subset K_{3}$ and a non-negative function $h$ with $h=-C_{1}C_{2}+1$ on $K_{2}$ and $h=0$ on $\mathbb{R}^{2n+1}\setminus K_{3}$. Let $(\varphi_{t})_{t\in[0,1]}$ be the contact flow generated by the contact Hamiltonian $h$. Then by Theorem 1.1 there exists a compactly supported null path $(\psi_{t})_{t\in[0,1]}$ connecting the identity and $\varphi_{-1}$. We pick the sets $K_{2}$ and $K_{3}$ such that $\varphi_{t}(K_{1})\subset K_{2}$ for all $t\in[-1,1]$. Hence on $K_{1}$ the flow $(\varphi_{t})_{t\in[0,1]}$ coincides with a reparametrization of the standard Reeb flow. Moreover, from the expressions (2)-(4) for the null path in this case, one sees that if we pick $K_{2}$ and $K_{3}$ large enough we can assume that $\varphi_{t}(\psi_{t}(K_{1}))\subset K_{2}$ for all $t\in[0,1]$. Then $(g_{t})_{t\in[0,1]}$ with $g_{t}=\varphi_{t}\circ\psi_{t}$ is a positive loop based at $\mathrm{id}$ with $\alpha\left(X_{t}^{g}\right)>-C_{1}C_{2}$ on $K_{1}$.
@@
Let $f\in\mathcal{D}_{c}(M)$ be compactly supported in some compact set $K$ let $U_{1},\cdots,U_{m}$ be a finite cover of $M$ by Darboux charts. The fragmentation Lemma [5] implies that there exist $f_{1},\cdots,f_{j}$ such that $f=f_{1}\circ\cdots\circ f_{j}$ and $f_{i}\subset\mathcal{D}_{c}(U_{k(i)})$. It follows from Theorem 1.1 that there exist null paths from $\mathrm{id}$ to $f_{i}$. Note that the paths constructed in the proof of Theorem 1.1 are compactly supported in $U_{k(i)}$ if $f_{i}$ is compactly supported. By concatenating and right translating these paths we obtain a compactly supported null path from the identity to $f$.
@@
By Proposition 3.1 for every fixed neighbourhood $\mathcal{W}_{i}\subset\mathcal{D}_{c}(U_{l(i)})\cap\mathcal{U}$ and each $t\in[0,1]$ there exists a neighbourhood $\mathcal{V}_{t}$ of $g_{t}$ in $\mathcal{D}_{c}(U_{l(i)})$ such that each $h\in\mathcal{V}_{t}$ is connected to $g_{t}$ by a null path contained in $\mathcal{W}_{i}$. Pick a subdivision $10=t_{0}\leq\cdots\leq t_{m}=$ 1 such that $g_{t_{j+1}}\in\mathcal{V}_{t_{j}}$ for all $0\leq j\leq m$. Then $g_{t_{j}}$ and $g_{t_{j+1}}$ are connected by a null path contained in $\mathcal{W}_{i}$. The concatenation of these paths is a null path between $\mathrm{id}$ and $g_{t}^{i}$ in $\mathcal{W}_{i}$.
@@
In case $M$ is closed, the all the topologies described above coincide, in particular the Whitney $C^{\infty}$-topology is locally contractible on $\mathcal{D}(M)$. As before let $\varphi_{t}$ be the Reeb flow of some contact form. Left composing each of the null paths obtained above with $(\varphi_{\epsilon t})_{t\in[0,1]}$ we obtain positive paths between $f_{t_{i}}$ and $f_{t_{i+1}}$ contained in $\mathcal{U}_{k(i)}$. Their concatenation is a positive path homotopic to $(f_{t})_{t\in[0,1]}$.

### v2

@@ 1. Introduction
Let $(M,\xi)$ be a (not necessarily closed) cooriented contact manifold of dimension $2n+1$ for $n\geq 1$. This means that the hyperplane distribution $\xi\subset TM$ is globally the kernel of a $1$-form $\alpha$ such that $\alpha\wedge(d\alpha)^{n}$ is a volume form on $M$.
There are two natural groups associated with $M$: the group of diffeomorphisms $\operatorname*{Diff}(M)$ and the group of contactomorphisms $\operatorname*{Cont}(M,\xi)$, i.e., the subgroup of diffeomorphisms preserving the contact structure $\xi$. In this paper we will mostly consider the following subgroups of $\operatorname*{Diff}(M)$: the group $\operatorname*{Diff}_{0}(M)$ of diffeomorphisms isotopic to the identity, the group $\operatorname*{Diff}_{c}(M)$ of compactly supported diffeomorphisms isotopic to the identity via compactly supported diffeomorphisms and their contact analogues $\operatorname*{Cont}_{0}(M,\xi)$ and $\operatorname*{Cont}_{c}(M,\xi)$.
@@ 1.1. Positivity for contactomorphisms and diffeomorphisms
In [12] the authors introduce the notion of a positive path of contactomorphisms (see also [7]) as follows. Any smooth path $(\varphi_{t})_{t\in I}$ in $\operatorname*{Cont}(M,\xi)$ defines a time-dependent contact vector field $(X_{t}^{\varphi})_{t\in I}$ on $M$. A path is called positive (non-negative) if $\alpha\left(X_{t}^{\varphi}\right)>0$ ($\geq 0$) for all $t$, where $\alpha$ is any contact form defining the chosen coorientation of $\xi$. The notion of positivity can be used to define a bi-invariant relation on $\operatorname*{Cont}(M,\xi)$ (and its universal cover $\widetilde{\operatorname*{Cont}(M,\xi)}$), see e.g. [12]. For closed contact manifolds this relation turns out to be a partial order if and only if there are no positive (contractible) loops of contactomorphisms [12].
The notion of positivity has turned out to be connected to several other known phenomena in contact geometry, such as for example the existence of closed Reeb orbits [2], contact (non-)squeezing [11, 4], the existence of invariant norms [10, 14], translated points [27, 2, 5, 19], the existence of quasimorphisms [16, 17], (non-) displaceability of pre-Lagrangians [8, 23], loose Legendrians [22] or supporting open book decompositions [19].
In principle the same notion of positivity can be defined for paths in $\operatorname*{Diff}(M)$, see [13, Remark 5.4]. Since any path $(f_{t})_{t\in I}$ defines a time-dependent vector field $(X_{t}^{f})_{t\in I}$ on $M$, we can analogously define a path of diffeomorphisms to be positive (non-negative) if $\alpha\left(X_{t}^{f}\right)>0$ ($\geq 0$) for all $t$. Note however that this notion of positivity is in general not invariant under conjugation. Another major difference to the contact case is that there exist non-constant paths of diffeomorphisms with $\alpha\left(X_{t}^{f}\right)=0$. We call a path $(f_{t})_{t\in I}$ with this property a null path.
In this paper we show that the notion of positivity is trivial on the identity component of the diffeomorphism group $\operatorname*{Diff}_{0}(M)$ in the sense of the following theorems.
@@
Consider $\mathbb{R}^{2n+1}$ equipped with its standard contact structure. Every $f\in\operatorname*{Diff}_{0}(\mathbb{R}^{2n+1})$ is connected to $\mathrm{id}$ by a null path and a positive path of diffeomorphisms. If $f\in\operatorname*{Diff}_{c}(\mathbb{R}^{2n+1})$, it is connected to $\mathrm{id}$ by a compactly supported null path and a positive path that is arbitrarily close to $\mathrm{id}$ in the $C^{\infty}$-compact open topology outside of a compact set.
@@
The following shows that in $\mathbb{R}^{2n+1}$ it is possible to deform a given path of contactomorphisms that is positive outside a compact subset to an everywhere positive path of diffeomorphisms with the same endpoints.
@@
Consider $\mathbb{R}^{2n+1}$ equipped with its standard contact structure. Let $(f_{t})_{t\in[0,1]}$ be a path of contactomorphisms that is positive on the complement of some compact subset $K_{1}$. Then there exists a positive path of diffeomorphisms $(g_{t})_{t\in[0,1]}$ connecting $f_{0}$ and $f_{1}$ such that $g_{t}$ coincides with $f_{t}$ on the complement of some compact subset $K_{2}\supset K_{1}$.
For more general contact manifolds one can show the following.
@@
For the case of compactly supported diffeomorphisms the local contractibility of the $C^{\infty}$-topology implies the following.
@@
On many closed contact manifolds it is known that a contractible positive loop of contactomorphisms can never be contracted in $\operatorname*{Cont}_{0}(M,\xi)$ via non-negative loops, see e.g. [11, Theorem 1.11] for the case of $S^{2n+1}$. This changes for diffeomorphisms.
@@ 1.2. Positive paths between submanifolds and thermodynamic processes
As pointed out in [13, Remark 5.4] one can also define a notion of positivity for smooth families of embeddings into a contact manifold. Let $N$ be a smooth manifold and $\iota\colon N\times[0,1]\hookrightarrow M$ be a smooth family of embeddings into a cooriented contact manifold $(M,\xi=\ker\alpha)$. We assume that all embeddings are proper in the sense that the intersection of the image with any compact subset in $M$ is compact. Call $\iota$ positive/ non-negative/ null if $\alpha(d\iota(\partial_{t}))$ is positive/ non-negative/ null for any contact form $\alpha$ inducing the given coorientation, i.e., if the path $t\mapsto\iota(p,t)$ is positively/non-negatively transverse or tangent to $\xi$ for any $p\in N$.
@@
In [13] the authors look at positive (non-negative) isotopies of Legendrians in the context of thermodynamic processes. More precisely, they look at sets of equilibrium states in a thermodynamic phase space $\mathcal{T}$. The space $\mathcal{T}$ naturally carries a contact structure and is contactomorphic to the $1$-jet bundle $J^{1}\mathbb{R}^{n}\cong\mathbb{R}^{2n+1}$ of $\mathbb{R}^{n}$ (or more generally of any smooth manifold $X$) equipped with its canonical contact structure. Sets of equilibrium states naturally form Legendrian submanifolds in this space, i.e., submanifolds $L$ with $TL\subset\xi$. As before we assume that $\mathcal{T}$ can be covered by finitely many Darboux charts. Slow thermodynamic processes can be described in this language by compactly supported non-negative isotopies of Legendrians, see [13, Section 5]. In [13, Remark 5.4] the authors consider smooth (possibly non-Legendrian) isotopies between two Legendrians and raise the question if the non-negativity assumption is restrictive in this setting. The following corollary of Theorem 1.1 and Theorem 1.3 provides a negative answer.
@@
Let $L_{1},L_{2}$ be Legendrians in $\mathcal{T}$ that are isotopic via a compactly supported smooth isotopy. Then there exist a positive and a compactly supported null isotopy between $L_{1}$ and $L_{2}$. If $\mathcal{T}\cong J^{1}\mathbb{R}^{n}$ the statement is true for any two Legendrians $L_{1},L_{2}$ such that there exist $f\in\operatorname*{Diff}_{0}(J^{1}\mathbb{R}^{n})$ with $f(L_{1})=L_{2}$. Moreover, in this case the positive isotopy can be chosen arbitrarily close in the $C^{\infty}$-compact open topology to the constant isotopy in the complement of a compact subset.
@@ 1.3. Norms on groups of diffeomorphisms
A further direct consequence of Theorem 1.3 and Corollary 1.6 is that, in contrast to contact geometry [28, 26], the notion of contact Hamiltonian can not be used to define any meaningful Hofer-type (psudo-) norms or metrics on the group $\operatorname*{Diff}_{c}(M)$ or the isotopy class of any submanifold.
For instance given a path of diffeomorphisms $(f_{t})_{t\in[0,1]}$ on a closed manifold $M$ with generating time-dependent vector field $X_{t}$, one can define its Hofer length with respect to a contact form $\alpha$ as
@@ Acknowledgements
I thank Leonid Polterovich for his comments on a first version of the paper and for encouraging to prove parts of Theorem 1.1 and Theorem 1.2. I also thank Egor Shelukhin, Souheib Allout and the anonymous referee for valuable feedback. This work was supported by a CRM-ISM and CIRGET postdoctoral fellowship, by the Fondation Courtois and by the Radboud Excellence Initiative.
@@ 2.1. Notations
Throughout this note $(M,\xi)$ will be a (not necessarily closed) contact manifold of dimension $2n+1$ for some $n\geq 1$. We assume that $(M,\xi)$ is cooriented, i.e., that $\xi$ is the kernel of some global $1$-form $\alpha$ such that $\alpha\wedge d\alpha^{n}$ is a volume form on $M$.
@@
A smooth path of diffeomorphisms $(f_{t})_{t\in I}$ is a family of diffeomorphisms which is smooth as a map $I\times M\rightarrow M$. Given a possibly time dependent vector field $X_{t}$, we will denote its flow at time $t$ by $\varphi_{t}^{X}$ and the time dependent vector field generated by a path of diffeomorphisms $f_{t}$ by $X_{t}^{f}$.
In the proofs we will often use the following combination of concatenation and right translation of two paths of diffeomorphisms starting at the identity to obtain a path from the identity to the composition of their endpoints. Given two paths of diffeomorphisms $(f_{t})_{t\in[0,1]}$ and $(g_{t})_{t\in[0,1]}$ starting at the identity one can obtain a smooth path $(h_{t})_{t\in[0,1]}$ from the identity to $f_{1}g_{1}$ as follows. Let $\tau\colon[0,1]\rightarrow\ [0,1]$ be an increasing smooth function with $\tau(0)=0$, $\tau(1)=1$ and $\mathrm{supp}(\tau^{\prime})\subset[1/2,1]$ and let $\eta\colon[0,1]\rightarrow\ [0,1]$ be an increasing smooth function with $\eta(0)=0$, $\eta(1)=1$ and $\mathrm{supp}(\eta^{\prime})\subset[0,1/2]$. Then the path $(h_{t})_{t\in[0,1]}=(f_{\tau(t)}g_{\eta(t)})_{t\in[0,1]}$ is smooth and generates the time dependent vector field
$X_{t}^{h}=\tau^{\prime}(t)X_{\tau(t)}^{f}\circ g_{\eta(t)}+\eta^{\prime}(t)df_{\tau(t)}\left(X_{\eta(t)}^{g}\right).$
Consequently, if $(f_{t})_{t\in[0,1]}$ and $(g_{t})_{t\in[0,1]}$ are both null paths, then so is $(h_{t})_{t\in[0,1]}$ by the definition of the functions $\tau$ and $\eta$. Note that this is in general not true for the composition $(f_{t}g_{t})_{t\in[0,1]}$ unless $f_{t}$ preserves the contact structure.
@@ 3.1. Diffeomorphisms of Darboux-balls
The proof of the following theorem uses a variation of Proposition 2.2 for the case of the standard contact structure. The main difference here is that for the standard contact structure the above vector fields used in Proposition 2.2 can be chosen linearly independent on all of $\mathbb{R}^{2n+1}$. This allows to extend the result to non-compactly supported diffeomorphisms.
@@
(1) $f=\varphi^{X}_{\tau_{1}\circ\Phi_{2}^{-1}}\circ\varphi^{Y}_{\tau_{2}\circ\Phi_{1}^{-1}}\circ\varphi^{Z}_{\tau_{3}}=\varphi^{X}_{\tau_{1}\circ\Phi_{2}^{-1}}\circ\varphi^{Y}_{\tau_{2}\circ\Phi_{1}^{-1}}\circ\varphi^{X}_{\epsilon}\circ\varphi^{Y}_{\tau_{3}\circ\varphi_{\epsilon}^{X}}\circ(\varphi^{X}_{\epsilon})^{-1}$
for $f\in\mathcal{W}_{\epsilon}$.
@@
Let $\varphi_{t}$ be the Reeb-flow of some contact form $\alpha$, i.e., $\varphi_{t}$ is the flow of a vector field $R$ uniquely defined by the equations $\alpha(R)=1$ and $d\alpha(R,\cdot)=0$. Then $\varphi_{t}$ is a positive path of diffeomorphisms preserving the contact form $\alpha$. Pick a null path of diffeomorphisms $(g_{t})_{t\in[0,1]}$ with $g_{0}=\mathrm{id}$ and $g_{1}=\varphi_{-1}f$. The composition $(\varphi_{t}g_{t})_{t\in[0,1]}$ is a positive path of diffeomorphisms from $\mathrm{id}$ to $f$.
@@
(4) $\varphi^{Y}_{s(\tau_{3}\circ\varphi_{\epsilon}^{X})}$ $=\left(x,y+\frac{st}{\epsilon},z+(\epsilon-x)\frac{st}{\epsilon}\right).$
@@
Let $\mathcal{U}_{co}$ be a neighbourhood of the identity in the $C^{\infty}$-compact open topology. For $|t|$ sufficiently small the Reeb flow $\varphi_{t}$ at time $t$ lies in $\mathcal{U}_{co}$ and from the above expressions one sees that, picking $|t|$ small enough, $\varphi_{t}$ is connected to $\mathrm{id}$ by a null path contained in $\mathcal{U}_{co}$. Given a compactly supported path $(f_{s})_{s\in[0,1]}$ and fixed $t$ small enough one gets a null path $g_{s}$ connecting $\mathrm{id}$ and $\varphi_{-t}\circ f_{1}$ that is arbitrarily close in the compact open topology to the identity outside of some compact subset. As before we get the desired positive path by looking at the composition $\varphi_{st}\circ g_{s}$.
@@
Let $(f_{t})_{t\in[0,1]}$ be a path of contactomorphisms that is positive outside a compact set $K_{0}$. We can without restriction assume that $f_{0}=\mathrm{id}$. Pick a compact set $K_{1}$ such that $f^{-1}_{t}(K_{0})\subset K_{1}$ for all $t\in[0,1]$. Given a contact form $\alpha$, let $C_{1}$ be the minimum of the time-dependent contact Hamiltonian inducing the path $(f_{t})_{t\in[0,1]}$. Note that $f_{t}^{\ast}\alpha=\rho_{t}\alpha$ for some family of positive functions $(\rho_{t})_{t\in[0,1]}$. Let $C_{2}$ be the maximum of the family $(\rho_{t})_{t\in[0,1]}$ on $K_{1}$. Pick compact subsets $K_{2},K_{3}$ with $K_{1}\subset K_{2}\subset K_{3}$ and a non-negative function $h$ with $h=-C_{1}C_{2}+1$ on $K_{2}$ and $h=0$ on $\mathbb{R}^{2n+1}\setminus K_{3}$. Let $(\varphi_{t})_{t\in[0,1]}$ be the contact flow generated by the contact Hamiltonian $h$. Then by Theorem 1.1 there exists a compactly supported null path $(\psi_{t})_{t\in[0,1]}$ connecting the identity and $\varphi_{-1}$. We pick the sets $K_{2}$ and $K_{3}$ such that $\varphi_{t}(K_{1})\subset K_{2}$ for all $t\in[-1,1]$. Hence on $K_{2}$ the flow $(\varphi_{t})_{t\in[0,1]}$ coincides with a reparametrization of the standard Reeb flow. Moreover, from the expressions (2)-(4) for the null path in this case, one sees that if we pick $K_{2}$ and $K_{3}$ large enough we can assume that $\varphi_{t}(\psi_{t}(K_{1}))\subset K_{2}$ for all $t\in[0,1]$. Then $(g_{t})_{t\in[0,1]}$ with $g_{t}=\varphi_{t}\circ\psi_{t}$ is a positive loop based at $\mathrm{id}$ with $\alpha\left(X_{t}^{g}\right)>-C_{1}C_{2}$ on $K_{1}$.
@@
Let $f\in\mathcal{D}_{c}(M)$ be compactly supported in some compact set $K$ let $U_{1},\cdots,U_{m}$ be a finite cover of $M$ by Darboux charts. The fragmentation Lemma [6] implies that there exist $f_{1},\cdots,f_{j}$ such that $f=f_{1}\circ\cdots\circ f_{j}$ and $f_{i}\in\mathcal{D}_{c}(U_{k(i)})$. It follows from Theorem 1.1 that there exist null paths from $\mathrm{id}$ to $f_{i}$. Note that the paths constructed in the proof of Theorem 1.1 are compactly supported in $U_{k(i)}$ if $f_{i}$ is compactly supported. By concatenating and right translating these paths we obtain a compactly supported null path from the identity to $f$.
@@
By Proposition 3.1 for every fixed neighbourhood $\mathcal{W}_{i}\subset\mathcal{D}_{c}(U_{l(i)})\cap\mathcal{U}$ and each $t\in[0,1]$ there exists a neighbourhood $\mathcal{V}_{t}$ of $g_{t}$ in $\mathcal{D}_{c}(U_{l(i)})$ such that each $h\in\mathcal{V}_{t}$ is connected to $g_{t}$ by a null path contained in $\mathcal{W}_{i}$. Pick a subdivision $0=t_{0}\leq\cdots\leq t_{m}=1$ such that $g_{t_{j+1}}\in\mathcal{V}_{t_{j}}$ for all $0\leq j\leq m$. Then $g_{t_{j}}$ and $g_{t_{j+1}}$ are connected by a null path contained in $\mathcal{W}_{i}$. The concatenation of these paths is a null path between $\mathrm{id}$ and $g_{t}^{i}$ in $\mathcal{W}_{i}$.
@@
In case $M$ is closed, all the topologies described above coincide, in particular the Whitney $C^{\infty}$-topology is locally contractible on $\mathcal{D}(M)$. As before let $\varphi_{t}$ be the Reeb flow of some contact form. Left composing each of the null paths obtained above with $(\varphi_{\epsilon t})_{t\in[0,1]}$ we obtain positive paths between $f_{t_{i}}$ and $f_{t_{i+1}}$ contained in $\mathcal{U}_{k(i)}$. Their concatenation is a positive path homotopic to $(f_{t})_{t\in[0,1]}$.